OpenStudy (anonymous):

|6x-7| = 6- |x-3| since it's in abs value, would it be -7-3 or 7+3? <-- when I add over |x-3|

OpenStudy (anonymous):

7+3 as absolute value do not take the minus sign

OpenStudy (anonymous):

so I would end up with 7x+10=6 instead of 7x-10=6?

OpenStudy (lgbasallote):

isn't it supposed to be 6x - 7 + x - 3 = -6?

OpenStudy (anonymous):

that's what i was thinking too

OpenStudy (anonymous):

where did the -6 come from?

OpenStudy (lgbasallote):

you make the right side negative

OpenStudy (lgbasallote):

to make the absolute value disappear*

OpenStudy (anonymous):

!!

OpenStudy (anonymous):

so sorry i made it worse >.<"

OpenStudy (anonymous):

what... how did you get -6? don't I just add -|x+3| over?

OpenStudy (lgbasallote):

yes |6x - 7| + |x-3| = 6 then, to make the absolute values disappear, you multiply the right side by -1

OpenStudy (anonymous):

i understand now

OpenStudy (lgbasallote):

i assume you're not familiar with that method @lookitswill ?

OpenStudy (anonymous):

no i'm not